Mr. Jacob is 55 yr old and Tom is 7 yrs old. In how many yrs will Mr. Jacob be 4 times as old as Tom?
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Mr. Jacob is 55 yr old and Tom is 7 yrs old. In how many yrs will Mr. Jacob be 4 times as old as Tom?
----------
Equation:
55+x = 4(7+x)
----
55+x = 28 + 4x
3x = 27
x = 9 years
Ans:: In 9 years
--------------------
Cheers,
Stan H.
--------------------

Answer by ewatrrr(24785)

x the number of years before Mr. Jacob will be 4 times as old as Tom
55+x = 4(4+x)
39 = 3x
13years = x, the number of years before Mr. Jacob will be 4 times as old as Tom
Checking...
68 = 4(17)